OpenStudy (anonymous):

which of the following expresses the coordinates of the foci of the comic section shown below? (x+2)^2 / 64+ (y-1)^2 / 81 = 1 its a circle

OpenStudy (anonymous):

\[\frac{ (x+2)^2 }{ 64} + \frac{ (y-1)^2 }{ 81 } = 1\]

OpenStudy (jdoe0001):

well... a circle is just an ellipse with the "a" and "b" components, or the minor and major axis equalling each other

OpenStudy (anonymous):

mmmhhh okkk

OpenStudy (anonymous):

sooo 8 and 9 then what would i do with this ??

OpenStudy (jdoe0001):

so the foci will be at the "c" distance from the center of the ellipse notice the BIGGER number is the 81 and that one is under the fraction with the "y" variable meaning the ellipse is a TALL ellipse, that is the major axis is over the y-axis  and the "c" distance for an ellipse is at \(\bf c=\sqrt{a^2-b^2}\qquad \cfrac{(x-{\color{blue}{ h}})^2}{{\color{brown}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{brown}{ a}}^2}=1\qquad foci\to ({\color{blue}{ h, k\pm}}c)\)

OpenStudy (anonymous):

ok soooo \[c=\sqrt{17}\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Foci is \[(-2,1\pm \sqrt{17}\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay omgoshhh thank you for showing me this n_n ! @jdoe0001

OpenStudy (jdoe0001):

yw